A five foot tall tree casts an 8 foot shadow on the ground. How far is it from the end of the shadow to the top of the tree?
Answer by stanbon(75887)

A five foot tall tree casts an 8 foot shadow on the ground. How far is it from the end of the shadow to the top of the tree?
----------
Use Pytagoras:
h^2 = 5^2 + 8^2
---
h^2 = 25 + 64
---
h^2 = 89
---
hypotenuse = sqrt(89) = 9.43 ft
================================
Cheers,
Stan H.